Home‎ > ‎Crew 160‎ > ‎

0108-DailySummaryReport

MDRS Daily Summary Report for sol 13 1/08/2016

Mars against Humanity

Executive Commander Reporting

Mission Status: Today the crew repaired and cleaned the Hab for the next crew rotation. Worked on reports and enjoyed more Star Wars.

Sol Activity Summary: A lazy start for everyone today as the crew rolled out of bed to start being busy bees in the Hab. A few crew members cleaned up the kitchen in the Hab and a few other crew members focused fixing the stair ways in the Hab. After this exciting endeavor, the crew went crazy…light sabers everywhere…just kidding. The crew focused on daily reports and enjoyed the rest of the day with Star Wars and more nerd fests about how Star Trek is better (intercepted now by Crew Commander – Star Wars is better).

Look Ahead Plan: Prepping for new crew arrival.

Anomalies in work: There are no anomalies to report.

Weather: The Martian sky was clear as apple pie today.

Crew Physical Status: Each crew member is doing fine!

EVA: None

Reports to be filed:
Final Commanders Report
Journalist Report
GreenHab Report
Engineering Report
End of Mission Statement

Support Requested: N/A